OpenStudy (anonymous):

Solve for Y showing all steps -10+2=5y-1

OpenStudy (zale101):

add 1 to both sides to leave 5y by itself

OpenStudy (anonymous):

SO I would add 1 to -10, 1, and 2? @Zale101

OpenStudy (zale101):

-10+2=5y-1 -10+2+1=5y-1+1

OpenStudy (zale101):

do your additions and subtractions now, and clear all that out

OpenStudy (anonymous):

I got y=-13/5 Would that be right? @Zale101

OpenStudy (zale101):

that's not right

OpenStudy (zale101):

-10+2+1=5y -10+3=5y -7=5y divide 5 to both sides y=-7/5

OpenStudy (zale101):

pay attention to the order of operations PEMDAS

OpenStudy (zale101):

do you know them? :)

OpenStudy (anonymous):

Thank you for your help, I guess I was just adding wrong :)